OpenStudy (anonymous):

Please explain:) Factor out the greatest common factor from the expression: 33x^7-18x^4

jimthompson5910 (jim_thompson5910):

the GCF of 33 and 18 is 3 (since this is the largest number that goes into both) the GCF of x^7 and x^4 is x^4 (again, using the same reasoning) So the GCF of 33x^7 and 18x^4 is 3x^4 Now factor out the GCF to get \[\Large 3x^4(11x^3-6)\]

jimthompson5910 (jim_thompson5910):

Notice how distributing and multiplying gives us the original expression again

OpenStudy (anonymous):

the GCF of 33 and 18 is 3....GCF of x^7 and x^4 is x^4 cause they r to the powers....GCF of 33x^7 and 18x^4 is 3x^4 and then you factor and the expression given by jim is correct =]

OpenStudy (anonymous):

okay so how would I go by factoring out the GCF to get: 3x^4(11x^3-6)

jimthompson5910 (jim_thompson5910):

first, write 3x^4 outside a set of parenthesis like so 3x^4( ) Now divide 33x^7 by 3x^4 to get (33x^7)/(3x^4) = (33/3)x^(7-4) = 11x^3 This is the first term that will go in the parenthesis Now divide -18x^4 by 3x^4 to get (-18x^4)/(3x^4) = (-18/3)x^(4-4) = -6 This is the second term that will go in the parenthesis So we now have 3x^4(11x^3 - 6)

OpenStudy (anonymous):

oh wow thank you Jim, I need a tutor so bad for this factoring thing but you showed my how so maybe it will be easier for me for the next problem:)

OpenStudy (anonymous):

you are simplifying the original equation...so you find whats common and then you take it out side of the ( ) ....which is 3x^4( ) now you put in 11x^3 in the ( ) b.c the product of 3x^4(11x^3) is 33x^7...and then do 3x^4( -6) = -18x^4....and since the number outside...3x^4...is the same combine the two numbers in the parentheses....3x^4(11x^3 - 6) in order to check if your answer is right just follow distributive property =]

OpenStudy (anonymous):

HEY GUYS WHAT'S UP? I have a question and I need it explained step by step for my test, PLEASE HELP! The question is factor out the greatest common factor from the expression: 20x^8+30x^7-15x^4?

jimthompson5910 (jim_thompson5910):

just answered it back at your other post